OpenStudy (anonymous):

The sum of the solutions of (5x+3)^2=16 is?

jhonyy9 (jhonyy9):

5x+3=4 5x=1 x=1/5

OpenStudy (anonymous):

...sum of solutions?

OpenStudy (anonymous):

and 5x + 3 = -4 5x = -7 x = -7/5 sum is -6/5

OpenStudy (anonymous):

(5x+3)(5x+3)=16 25x^2+30x+9-16=0

OpenStudy (anonymous):

Don't forget the plus or minus!\[5x+3=\pm 4\]\[x = \dfrac{-3\pm4}{5}=\left\{\dfrac{-3 + 4}{5}, \dfrac{-3 - 4}{5}\right\}\]Add the two results together:\[\dfrac{-3 + 4}{5}+ \dfrac{-3 - 4}{5}=-\dfrac{6}{5}\]For more solutions to textbook problems, be sure to check out http://www.slader.com/s/eWFrZXlnbGVl. I submit a lot of content there as well.